The combined SAT scores for the students at a local high school are normally distributed with a mean of 1530 and a standard deviation of 296. The local college includes a minimum score of 820 in its admission requirements.

What percentage of students from this school earn scores that satisfy the admission requirement?
P(X > 820) =  %

Enter your answer as a percent accurate to 1 decimal place (do not enter the "%" sign). Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Solution :

Given that,

mean = = 1530

standard deviation = = 296

P (x > 820 )

= 1 - P (x < 820 )

= 1 - P ( x -  / ) < ( 820 - 1530 / 296)

= 1 - P ( z < -710 / 296 )

= 1 - P ( z < -2.399 )

Using z table

= 1 - 0.0082

= 0.9918

Probability = 0.9918 =99.18%